### v2

# Full-Duplex-Bench v1.5: Evaluating Overlap Handling for Full-Duplex Speech Models

Abstract
Full-duplex spoken dialogue systems promise to transform human-machine interaction from a rigid, turn-based protocol into a fluid, natural conversation. However, the central challenge to realizing this vision, managing overlapping speech, remains critically under-evaluated. We introduce Full-Duplex-Bench v1.5, the first fully automated benchmark designed to systematically probe how models behave during speech overlap. The benchmark simulates four representative overlap scenarios: user interruption, user backchannel, talking to others, and background speech. Our framework, compatible with open-source and commercial API-based models, provides a comprehensive suite of metrics analyzing categorical dialogue behaviors, stop and response latency, and prosodic adaptation. Benchmarking five state-of-the-art agents reveals two divergent strategies: a responsive approach prioritizing rapid response to user input, and a floor-holding approach that preserves conversational flow by filtering overlapping events. Our open-source framework enables practitioners to accelerate the development of robust full-duplex systems by providing the tools for reproducible evaluation [^1] Code and data are available at https://github.com/DanielLin94144/Full-Duplex-Bench.
Index Terms—  Full-Duplex Model, Spoken Dialogue System, Evaluation Benchmark
1 Introduction
Spoken dialogue systems are poised to evolve from command-and-response tools into genuine conversational partners. The cornerstone of this transition is full-duplex capability—the ability to speak and listen simultaneously—mirroring the dynamics of human conversation [1]. This concurrent processing enables fluid interactions, such as mid-utterance corrections and backchannels, which are essential for applications demanding high responsiveness, from in-car assistants to real-time translation.
Fig. 1: Overview of the evaluation framework for full-duplex speech models. User speech (top) overlaps with model output (bottom) in four controlled scenarios. We analyze the model’s post-overlap response across three dimensions: categorical behaviors, interaction timing, and adaptive speech features.
The defining characteristic of natural dialogue is not the absence of interference, but the ability to manage overlapping speech in a meaningful way. Overlap is not an edge case but a common conversational event, accounting for over 40% of turns [2, 3, 4]. It plays diverse roles, from user interruptions and backchannels to side conversations and ambient speech [5, 6]. A system that cannot handle overlap gracefully remains locked in a transactional, half-duplex paradigm, resulting in truncated responses, awkward silences, and degraded interaction quality.
Recent research has explored a variety of architectures to enable full-duplex behavior. Cascaded pipelines decompose the system into ASR, LLM, and TTS components and coordinate them with token-level control or time-sliced windows, as in FSM [7] and MiniCPM-Duplex/Duo [8, 9]. End-to-end approaches, such as dGSLM [10], SyncLLM [11], Moshi [12], and NTPP [13] learn to internalize turn-taking directly through joint speech modeling, while systems like SALMONN-omni [14], and MinMo [15] introduce state tokens or dialogue managers to refine floor control. Despite this progress, reproducibility remains a challenge: only a few models (e.g., Freeze-Omni and Moshi) release public checkpoints, while most state-of-the-art systems—including Gemini Live, GPT-4o Realtime, and Nova Sonic—are accessible only through closed APIs.
Evaluation methodology has lagged even further behind. Mainstream speech benchmarks [16, 17, 18, 19, 20] focus on single-turn, half-duplex settings and fail to capture the dynamics of overlap phenomena. Human evaluations offer nuanced judgments but are expensive and difficult to reproduce. Corpus-level analyses rely on pause statistics and floor-transfer offset [10], which scale well but obscure scenario-specific behavior and ignore semantic appropriateness. Classifier-based approaches such as Talking Turns [21] automate turn-change detection but remain tied to specific training corpora, limiting their generality.
To address these gaps, we present Full-Duplex-Bench v1.5, an extension of our earlier work [22], which offers the first systematic, automated benchmark for overlap handling. Our framework streams audio in real time to both open-weight and API-based systems, introduces four controlled overlap scenarios—Interruption, Backchannel, Talking to Others, and Background Speech—and evaluates not only what a model says but also when and how it responds, through metrics that jointly measure dialogue behavior, timing, and prosodic adaptation. We benchmark five state-of-the-art systems and reveal two contrasting strategies for overlap management, quantifying their trade-offs and providing the community with a reproducible testbed for developing robust full-duplex dialogue systems.
Fig. 2: Illustration of the four controlled overlap scenarios. User speech (top) and model speech (bottom) share a timeline: (1) Interruption: user barges in with a new request; (2) Backchannel: brief acknowledgment (uh-huh); (3) Talking to Other: user addresses someone else; (4) Background Speech: far-field third-party talk not meant for the model.
2 Full-Duplex-Bench-v1.5
2.1 Interaction Framework
Our framework simulates a live conversation by streaming pre-recorded user audio to a model while simultaneously capturing its output. This approach is model-agnostic, making no assumptions about internal architecture and thus supporting both open-source models and closed commercial APIs. The core evaluation primitive is a trial where a controlled overlap event is introduced while the model is speaking. The model’s audio output is automatically segmented into pre-overlap event and post-overlap event regions relative to the user’s speech, enabling precise, localized analysis of its reaction (See the timeline illustration in Fig.1). We adopt these names consistently throughout the paper; subsequent sections refer to them simply as the pre-overlap event and post-overlap event. The framework is fully extensible, allowing researchers to introduce new audio assets, languages, or custom scenarios.
2.2 Controlled Overlap Scenarios
We design four scenarios, illustrated in Fig. 2, to probe distinct and essential conversational capabilities.
2.2.1 User Interruption
Capability Tested. Reactive turn-yielding and semantic repair. When a user barges in (e.g., “Wait, I want to …"), an effective system must cede the floor rapidly and address the new query. This is critical in safety-conscious contexts like in-car navigation.
Data. We synthesize 200 contextually relevant interruptions using the same speaker voice as the initial query, with no acoustic channel differences.
2.2.2 User Backchannel
Capability Tested. Filtering non-floor-taking cues. Listeners often produce short affirmations (e.g., “uh-huh") to signal engagement, not to take the turn. The model must ignore these cues to maintain conversational flow.
Data. We synthesize 99 backchannel utterances (e.g., yeah, right, mm-hmm) from a curated list [23] using the same speaker voice.
2.2.3 User Talking to Others
Capability Tested. Addressee detection and social appropriateness. When a user addresses another person (e.g., “Hey Grandma, I’m fine. Don’t worry."), the model must recognize that it is not the intended recipient and gracefully resume. This is crucial for multi-party settings.
Data. We synthesize 100 utterances semantically directed to another person. To simulate the user speaking away from the device, we apply acoustic processing: volume reduced by 8 dB, a high-shelf filter attenuating frequencies above 4 kHz by 5 dB, and two reflections added at 45 ms (–6 dB) and 120 ms (–12 dB). This configuration approximates off-axis, far-field speech with early reflections, which is common in distant conversational setups [24].
2.2.4 Background Speech
Capability Tested. Robustness to ambient acoustic interference. Real-world environments contain incidental speech not directed at the system. The model must remain focused on its task and not be derailed by this noise.
Data. We synthesize 100 utterances with a different speaker voice on an unrelated topic. To simulate distant, environmental speech, we reduce the volume by 15 dB, apply a 3 kHz low-pass filter, and add echo with a 100 ms delay (–10 dB). The low-pass + attenuation + echo combination approximates distant, bandwidth-limited speech [25].
2.3 Evaluation Metrics
2.3.1 Dialogue Behavior
To understand a model’s semantic strategy, we classify its post-overlap event response into one of five categories using GPT-4o, informed by time-aligned transcripts from an ASR model (Parakeet-TDT) [26]:
  • Respond: Addresses the content of the overlapping speech.
  • Resume: Ignores the overlap and continues its prior utterance.
  • Uncertain: Expresses confusion (e.g., “Could you repeat that?”).
  • Unknown: Produces an irrelevant response or remains silent.
2.3.2 Interaction Timing
For each overlap event, we compute two timing measures (Fig. 1 illustrates these time intervals):
Stop latency is the interval from the onset of overlapping user speech to the moment the model stops speaking:
$t_{\text{stop}}=t_{\text{model\_stop}}-t_{\text{user\_start}}.$ (1)
Response latency is the interval from the end of the overlapping speech to the model’s next utterance:
$t_{\text{resp}}=t_{\text{model\_start}}-t_{\text{user\_end}}.$ (2)
Here, $t_{\text{user\_start}}$ and $t_{\text{user\_end}}$ mark the start and end of user speech, while $t_{\text{model\_stop}}$ and $t_{\text{model\_start}}$ mark when the model stops and resumes speaking. Both $t_{\text{stop}}$ and $t_{\text{resp}}$ are in seconds, with speech activity detected by Silero-VAD [27].
Table 1: Scenario-wise expected outcomes. Each row specifies the desired categorical behavior and latency profile for that overlap scenario.
Scenario Expected Behavior Stop Latency Response Latency
User Interruption Respond Low Low
User Backchannel Resume High Low
Talking to Others Resume High Low
Background Speech Resume High Low
2.3.3 Scenario-wise Expected Outcome
We summarize the expected post-overlap behaviors and the corresponding desirable stop and response latency profiles in Table 1.
In particular, Stop Latency serves as an indicator of overlap awareness, with its optimal value varying by scenario: it should be short when rapid yielding is required and may be longer when the system is expected to hold the floor. In contrast, Response Latency captures the post-overlap gap and is ideally minimized across all scenarios.
User Interruption. Expected behavior is Respond: the system should promptly yield and address the overlapped intent. Thus we target a high Respond rate with low Stop Latency and low Response Latency; continuing the prior turn or failing to react are primary errors.
User Backchannel. Expected behavior is Resume: In conversation analysis, generic backchannels (e.g., “uh-huh”, “yeah”) conventionally invite the speaker to continue rather than take the floor [28]. We target a high Resume rate and minimal responses to backchannels; Stop Latency should be as high as possible (the model does not readily halt), while Response Latency should remain low to avoid unnecessary gaps.
User Talking to Others & Background Speech. Expected behavior is Resume, with brief addressee checks (Uncertain) acceptable. We target high Resume (with some Uncertain) and near-zero responses to others; Stop Latency should high (holding the floor), whereas Response Latency should be low.
3 Evaluation Setup
We evaluate full‑duplex speech models via their streaming interfaces:
Table 2: Behavioral response distribution across overlap scenarios, with average stop and response latencies (s). Boldface indicates the best performance. Rows with a light-colored background highlight the desired behavior for each scenario.
Scenario Class / Metric Freeze-Omni Moshi Gemini Sonic GPT-4o
[…]
Table 3: Prosodic feature shifts (Pre $\rightarrow$ Post) during Respond trials for user_intr. Arrows next to feature names indicate the expected trend. Cells show actual change ($\Delta$) with $p$-values; significant increases are in green, decreases in red, and non-significant results are gray.
Feature Freeze-Omni Moshi Gemini Sonic GPT-4o
WPM $\uparrow$ +1.03 (.79) +59.55 (<.001) +18.09 (.037) -11.63 (.14) +19.04 (<.001)
Pitch Mean $\uparrow$ -1.02 (.55) -8.94 (<.001) +3.28 (.24) +0.99 (.69) +5.60 (.002)
Pitch SD $\uparrow$ -0.08 (.93) -9.54 (<.001) -0.78 (.73) +0.57 (.71) +2.79 (.024)
Intensity Mean $\downarrow$ -1.50 (.001) -4.02 (<.001) -0.40 (.29) -1.88 (.010) -0.07 (.75)
Intensity SD $\uparrow$ +6.06 (<.001) +4.81 (<.001) +0.07 (.80) +8.62 (<.001) -2.26 (<.001)
UTMOSv2 $\rightarrow$ -0.07 (.051) -0.10 (.12) -0.08 (.13) -0.21 (<.001) +0.06 (.09)
4 Results
4.1 Scenario-wise Outcomes
Table 2 reports behavioral distributions and latencies across four overlap scenarios. Rows with shaded backgrounds indicate the desired outcome for each case.
4.1.1 User Interruption
The goal is rapid yielding and immediate handling of the new intent (Respond with low $t_{\text{stop}}$$t_{\text{resp}}$). GPT-4o exhibits the strongest responsiveness (Respond=0.78) and extremely fast yielding ($t_{\text{stop}}{=}0.23$ s), with Freeze-Omni closely following (Respond=0.72) and achieving the shortest $t_{\text{resp}}$ (1.35 s). In contrast, Gemini and Sonic frequently continue their prior turn (Resume=0.55/0.71) and are slow to yield ($t_{\text{stop}}>2$ s), reflecting excessive floor-holding under true interruption. Moshi is intermediate in timing but leaves many user intents unaddressed (Respond=0.50).
4.1.2 User Backchannel
[…]
4.1.3 User Talking to Others
[…]
4.1.4 Background Speech
The system should resume speaking quickly after noise but avoid treating it as an interruption. Sonic best matches this target (Resume=0.98, Respond=0.01), though with conservative gap management ($t_{\text{resp}}=2.76$ s). In contrast, Gemini and Freeze-Omni often respond inappropriately (0.70/0.63), while GPT-4o nearly always yields (Respond=0.93) with near-zero stop latency, again over-accommodating. Moshi remains unstable (Unknown=0.71).
4.2 Cross-Model Patterns and Error Modes
A consistent trade-off emerges between rapid responsiveness and robust floor control. GPT-4o excels at true interruptions (fastest $t_{\text{stop}}$ and highest Respond), but systematically misfires under talking_other and bkg_speech. Sonic and Gemini exhibit strong addressee discrimination (high Resume) yet are too conservative under genuine interruptions. Freeze-Omni offers the most balanced profile but over-reacts to background speech. Moshi shows instability across all non–floor-taking settings (Unknown high).
Timing-wise, effective human-like repair typically requires $t_{\text{resp}}\!\leq\!1.5$ s. GPT-4o achieves such gaps but often with the wrong action (false Respond), whereas Sonic and Gemini leave long gaps (2.0–2.7 s) that may harm perceived conversational flow.
Takeaways. Best at interruptions: GPT-4o (fast yielding), Freeze-Omni (shortest $t_{\text{resp}}$). Best at filtering: Sonic (backch/bkg_speech) and Gemini (talking_other). Overall, systems optimized for fast yielding must strengthen addressee detection and backchannel filtering, while those optimized for robust floor holding must improve decisiveness under true interruptions.
5 Prosodic and Quality Shifts
Beyond categorical behaviors and timing, our benchmark also evaluates how models adapt their speech prosody and perceptual quality during overlap. Prosodic adaptation is a key marker of conversational competence: humans naturally modulate tempo, pitch, and energy when resuming after an interruption, signaling turn re-entry and preserving conversational naturalness.
Specifically, we investigate whether models modulate their delivery within an utterance when responding to a user interruption. For trials labeled respond, we compare the segment immediately before the overlap (pre-overlap event) with the repair segment after the overlap (post-overlap event). We evaluate speaking rate (WPM), pitch (mean, standard deviation (SD)), intensity (mean, standard deviation (SD)), and predicted MOS (UTMOSv2) using paired $t$-tests ($p<0.05$). Results are summarized in Table 3. To promptly respond to user interruptions, we expect slight increases in speaking rate (WPM) and a subtle pitch/energy lift at re-entry, without large intensity swings. MOS should not degrade; large intensity SD spikes or rate jumps indicate brittle control rather than skillful adaptation.
Comparing pre-overlap event to post-overlap event on respond trials (Table 3) reveals two dominant adaptation regimes. First, a tempo/pitch–lift regime in which GPT-4o (and, to a lesser extent, Gemini) systematically accelerates and elevates pitch—accompanied for GPT-4o by greater pitch variability—consistent with an emphatic, floor-reacquisition strategy. Second, a soft-but-dynamic intensity regime in which Freeze-Omni and Sonic reduce mean intensity while increasing intensity variability, yielding quieter yet more contoured restarts with minimal change in rate or pitch. Moshi departs from both patterns: it exhibits an extreme speed-up alongside decreases in pitch level and variability and a drop in mean intensity with higher intensity variability, suggesting hurried, less melodic, and brittle re-entry control. Predicted quality (UTMOSv2) is largely stable pre $\to$ post, with only a small decline observed for Sonic.
Takeaway. Strong models make well-controlled increases in tempo and pitch to signal turn re-entry while keeping speech natural. In contrast, overly fast or uneven intensity changes lead to rushed and unstable responses. These results suggest that future systems should adopt prosodic adjustments to maintain fluency and conversational appropriateness.
6 Conclusion
Managing overlap is a core competency for real-time conversational AI. We present Full-Duplex-Bench v1.5, a fully automated, scenario-controlled benchmark that makes overlap measurable by formalizing expected behavior and timing ($t_{\text{stop}},t_{\text{resp}}$) across four representative cases, enabling reproducible comparison beyond turn-based evaluation. Across five state-of-the-art systems, we find a stable trade-off between responsiveness and floor holding: fast yielders excel on true interruptions but over-accommodate incidental speech, while robust holders resist non-addressed input yet delay necessary repairs. Our metrics provide scenario-specific latency targets and behavior distributions that make these differences comparable. By open-sourcing tasks, metrics, and code, Full-Duplex-Bench v1.5 offers a practical yardstick to diagnose weaknesses, track progress beyond half-duplex paradigms, and engineer systems that handle the fluid dynamics of human conversation with greater fluency and social awareness.